OpenStudy (anonymous):

We have measured that a person who jumps up vertically from the ground up can reach a height of h = 70 cm. (a) Determine the starting velocity va which the person has at the moment he let the earth

OpenStudy (anonymous):

use the equations for constant acceleration: u = ?, s = 0.70 m, g = -9.8 cm/s^-2, v = 0 v^2 = u^2 - 2gs 0 = u^2 - 2 * 9.8 * 0.7 u^2 = 19.6 * 0.7 = 13.72 u = sqrt 13.72 = 3.7 m/s

OpenStudy (anonymous):

what is u?

OpenStudy (turingtest):

initial velocity